Find the capacitance of a series plate of capacitor:

  • C1 = 3.4x10^-3F
  • C2 = 5.0x10^-13F
  • C3 = 4.3x10^-9F
  • Vt = 500v
1
Expert's answer
2021-04-19T17:09:51-0400

The total capacitance of the series capacitors can be found as follows (see https://www.allaboutcircuits.com/textbook/direct-current/chpt-13/series-and-parallel-capacitors/):


"C = \dfrac{1}{\dfrac{1}{C1} + \dfrac{1}{C2}+\dfrac{1}{C3}}"

Substituting the values, obtain:


"C = \dfrac{1}{\dfrac{1}{3.4\times 10^{-3}} + \dfrac{1}{5\times 10^{-13}}+\dfrac{1}{4.3\times 10^{-9}}} \approx 4.999\times 10^{-13}F"

Answer. "4.999\times 10^{-13}F".